OpenStudy (anonymous):

MEDAL AND FAN!!!!!!!!! The hypotenuse of a right triangle is 20 centimeters, and the shorter leg is 12 centimeters. Find the length of the other leg. 2 cm 4 cm 16 cm

OpenStudy (anonymous):

you know pythagoras theorem?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then what difficulty you are facing here?

OpenStudy (anonymous):

wait wouldnt it be 16 cm?

OpenStudy (anonymous):

yes, absolutely!

OpenStudy (anonymous):

thnx :)

OpenStudy (anonymous):

welcome, u didn't need my help at all :)

OpenStudy (anonymous):

i have another question im stuck on

OpenStudy (anonymous):

SIMPLIFY 1/5 times square root of 5n^3 times square root of 20n^5

OpenStudy (anonymous):

1/5 outside of square root signs?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so, it will be 1/5 sqrt(5*20)sqrt(n^3*n^5)= 1/5* 10* n^4= 2n^4